OpenStudy (anonymous):

Given: ABC is similar to XYC. Enter answer as a decimal. If BY = 5, BC = 20, AC = 18, then XC =

OpenStudy (amistre64):

cant really determine a picture from this; you have a BY that might be a typo

OpenStudy (anonymous):

it is by

OpenStudy (amistre64):

is this a triangle inside another? with either AB or XY in the middle as a line segment?

OpenStudy (anonymous):

XY IS in the middle yes

OpenStudy (amistre64):

BA, is the left side of this then; B up top?

OpenStudy (anonymous):

yep correct

OpenStudy (anonymous):

that is right

OpenStudy (amistre64):

5 = 20 ------ n = 5(18)/20 n = 18 n = 18/4 20 - 18/4 = 80-18/4 = 62/4 .... is that an option :)

OpenStudy (amistre64):

15.5

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

nope its not right i dont know why :/

OpenStudy (amistre64):

i spose we coulda went an easier route and did: 15=20 ------ n = 18 n = 15(18)/20 = 3(18)/4 = 3(9)/2 = 27/2 = 13.5 .....

OpenStudy (amistre64):

i hit the 5 instead of the 3 ....

OpenStudy (anonymous):

corecct :D

OpenStudy (anonymous):

would u mind helping me with one more your the only one that seemd to be able to help me

OpenStudy (amistre64):

i can try....

OpenStudy (anonymous):

The sides of a polygon are 3, 5, 6, 8, and 10. The perimeter of a similar polygon is 40. Find the sides of the second polygon (from smallest to largest). Put answers in decimal form (hundredths).

OpenStudy (amistre64):

k(3+5+6+8+10) = 40 k = 40/32 = 20/16 = 10/8 = 5/4 the sides are: (2)(1) 3(5)/4 = 3.75 5(5)/4 = 6.25 6(5)/4 = 7.5 8(5)/4 = 10 10(5)/4 = 12.5 -------------- 40.00 adds up lol

OpenStudy (anonymous):

correct you dont know how much your helping e i been stuck on this stuff for like a week

OpenStudy (amistre64):

did you try yelling at it?

OpenStudy (anonymous):

haha yes

OpenStudy (anonymous):

Given: rectangle ABCD is similar to rectangle ZBXY. Enter answer as a simplified, improper fraction. If BX = 8, BZ = 3, DC = 5, then XC = now i tryed this one but i dont get what i did wrong arent you supposed to cross mulitply

OpenStudy (amistre64):

where do i put the letters?

OpenStudy (anonymous):

bxc at the top then zy under it then ad at the bottom

OpenStudy (amistre64):

good enough?

OpenStudy (amistre64):

my computer froze....

OpenStudy (amistre64):

a is to 8 ------- 5 is to 3

OpenStudy (amistre64):

a = 40/3 ; its a corss multiply yeah then 8 + 40/3 = 8(3) +40/3 = 24+40/3 XC = 64/3

OpenStudy (anonymous):

nope :/

OpenStudy (amistre64):

then the diagram aint right :)

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

Given: rectangle ABCD is similar to rectangle ZBXY. Enter answer as a simplified, improper fraction. If ZY = 5, XC = 3, DC = 4, then XY =

OpenStudy (anonymous):

thats all i got

OpenStudy (amistre64):

and how does that match up with the graph i posted of it?

OpenStudy (amistre64):

that is a different problem than th one you posted right before

OpenStudy (anonymous):

oh sorry